What is the smallest 4- digits number that is divisible by 9?
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

What is the smallest 4- digits number that is divisible by 9?
-----------------
1000/9 = 111.111...
--> 112*9 = 1008

Answer by Theo(13342)

the smallest 4 digit number would be 1000, i believe.
1000 / 9 = 111.1111
111 * 9 = 999
112 * 9 = 1008
i believe 1008 is the smallest 4 digit number that is divisible by 9.
that's my guess.